OpenStudy (anonymous):

find the slope of the secant line to the function f(x)= sqrt(x) + 2 between x=1 and x=16

OpenStudy (anonymous):

\[m=\frac{y_2-y_1}{x_2-x_1}\] \[y_2=f(16)=\sqrt{16}+2=4+2=6\] \[y_1=f(1)=\sqrt{1}+2=3\] \[\frac{6-3}{16-1}=\frac{3}{15}=\frac{1}{5}\]